H2 (g) + I2 (g) ⇌ 2HI (g)


At 400. K for the reaction above, Kp = 62.5.

a) 0.40 mol of HI are introduced into 1.0 L flask & heated to 400. K, calculate pressure HI (g) before reaction.

(b) Calculate pressure (each species when equilibrium established) in (a).

(c) H2 (g), I2 (g), and HI (g)introduced to flask at 400. K, at pressures 0.40 , 0.40 , 5.0 atm, respectively. Calculate the Q and state direction in which reaction will proceed to reach equilibrium. 

(d) At a different temperature, H2 (g), I2 (g), HI (g) are introduced to flask at 0.040 0.020, 0.090 atm, respectively. If partial pressure HI (g) decreases 50%, calculate pressure of each species when equilibrium is established. 

(e) Using part (d)Carefully draw three curves - one for each of the three gases. The curves must show how the pressure of each of the three gases changes as equilibrium is established. Label each curve with the formula of the gas. 


1
Expert's answer
2021-04-07T11:32:51-0400

"K_p=K_c(RT)^{\u2206n}"


"K_c = \frac{0.40}{1.0}=0.4"


"R= 0.08314"


"\u2206n = 2-(1+1)=0"


"K_p = 0.4(0.08314\u00d7400)^0=0.4"